What can I do to solve for an from just a vertex at a origin and with the directrix of y= -3/4?
Answer by lwsshak3(11628)

solve for an equation from just a vertex at a origin and with the directrix of y= -3/4?
***
Given directrix (-3/4) shows this is a parabola that opens downward:
Basic form of equation for such a parabola with vertex at (0,0):
x^2=-4py
axis of symmetry:x=0
p=3/4 (distance from vertex to directrix on the axis of symmetry)
4p=3
equation: x^2=-3y